The sum of two numbers is 109. One number is 25 more than three times the other number. Find the two numbers.
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39616)
Answer by ikleyn(52776)

.

            You can easily solve it mentally.


Diminish "one number" by 25, for a minute.


Then the sum of 109 will decrease by 25 and will become 109 - 25 = 84.


Now it is 4 times the "other number".


Hence, the "other number" is 84/4 = 21.


Then the "one number" is 109 - 21 = 88.


Answer.  The numbers are 88 and  21.